# FreeRide: Harvesting Bubbles in Pipeline Parallelism

### v1

@@ 4.3. Profiling Bubbles and Side Tasks
Bubbles. To know the shapes of bubbles, FreeRide runs DeepSpeed, monitors its estimated SM occupancy and GPU memory consumption through the PyTorch profiler (pytorch_profiler, ), and automatically measures each bubble’s duration and available GPU memory. Since the pipeline schedule determines bubbles, this offline profiling is done only once for each model and pipeline scheduling on the same hardware platform.
Side tasks. After the programmer implements the side task, FreeRide profiles it with the automated profiling tool for its performance characteristics of GPU memory consumption and per-step duration, which FreeRide uses for side task management and GPU resource limit. For side tasks implemented using the iterative interface, this procedure is fully automated. The profiling tool runs the side task, monitors its GPU memory consumption, and records the timestamps at the start and end of RunNextStep() state transition for the per-step duration. For side tasks implemented using the imperative interface, the tool profiles GPU memory consumption in the same way. However, since the side task is not implemented step-wise, the automated profiling tool does not measure the per-step duration.
@@ 4.4. Side Task Management
FreeRide’s side task management has two main roles. First, upon receiving a new side task, the side task manager assigns it to a suitable side task worker. Second, when the pipeline training system adds bubbles to the side task manager, the side task manager initiates the state transitions of side tasks (Figure 4(a)) through RPCs. This way, the side tasks are only served during bubbles and do not compete for GPU resources with the main pipeline training workload.
@@
Input: $Task$ new side task, $Workers$ workers’ metadata.
1 $MinNumTasks\leftarrow\infty$, $SelectedWorker\leftarrow None$;
2 foreach $Worker\in Workers$ do
3     if $Worker.GPUMem>Task.GPUMem$ then
4          $NumTasks\leftarrow w.GetTaskNum()$;
5          if $NumTasks<MinNumTasks$ then
6               $MinNumTasks\leftarrow NumTasks$;
7               $SelectedWorker\leftarrow Worker$;
8
9if $SelectedWorker\neq None$ then
10     $SelectedWorker.Add(Task)$;
11
12 else
13     $RejectSideTask()$;
14
@@ 13     $RejectSideTask()$;
Algorithm 1 describes how the side task manager assigns side tasks to workers. When the side task manager receives a new side task together with its GPU memory requirement (through profiling, Section 4.3), it first filters out all workers with enough available GPU memory (lines 2—3). Then, from these workers, it selects the one with the smallest number of tasks (lines 4—7). If the side task manager has selected a worker, it will assign the side task to that worker (lines 8—9). Otherwise, it will reject the side task because of insufficient GPU memory (line 11).
Input: $Workers$ workers’ metadata.
1 while $SideTaskManagerIsRunning$ do
2     foreach $Worker\in Workers$ do
3          if $Worker.CurrentBubble\neq None$ then
4               if $Worker.CurrentBubble.HasEnded()$ then
5                    if $Worker.CurrentTask\neq None$ then
6                         $Worker.CurrentTask.PauseSideTask()$;
7
8                   $Worker.CurrentBubble\leftarrow None$;
9
10         if $Worker.HasNewBubble()$ then
11               $Worker.UpdateCurrentBubble()$;
12               if $Worker.CurrentTask=None$ then
13                    if $Worker.TaskQueue.IsEmpty()$ then
14                         $continue$;
15
16                   $Worker.CurrentTask\leftarrow Worker.TaskQueue.Next()$;
17              if $Worker.CurrentTask.IsCreated()$ then
18                    $Worker.CurrentTask.InitSideTask()$;
19
20              else if $Worker.CurrentTask.IsPaused()$ then
21                    $W.CurrentTask.StartSideTask()$;
22
23
@@ 14                         $continue$;
Algorithm 2 describes how the side task manager manages bubbles and side tasks during pipeline training. The side task manager iterates through all workers (line 2). If CurrentBubble has just ended for a worker, the side task manager will pause CurrentTask of the worker and clear CurrentBubble (lines 3—7). Upon a new bubble, the side task manager updates the CurrentBubble of this worker (lines 8—9). It then checks if the worker has a CurrentTask. If not, it will select the one with the smallest submission timestamp from TaskQueue as CurrentTask (lines 10–13). After that, the side task manager initiates InitSideTask() if the newly added CurrentTask is in CREATED state (lines 14—15); otherwise, its state is PAUSED and the side task manager initiates StartSideTask() (lines 16—17).
@@ 4.5. GPU Resource Limit
GPU Memory. FreeRide leverages MPS to impose GPU memory limit on side tasks, i.e., when a worker creates a side task, it sets GPU memory limits using MPS. The side task process triggers an out-of-memory (OOM) error when its memory consumption exceeds the limit, but other processes remain unaffected. However, FreeRide is also compatible with other mechanisms for limiting GPU memory, e.g., multi-instance GPU (MIG) (nvidia_developers_mig, ) or manually implemented accounting through intercepting CUDA kernel calls (strati_orion_2024, ).
@@ 4.6. Implementation
We use DeepSpeed 0.12.2 (deepspeed_0_12_2, ) as the framework for pipeline training. We modify DeepSpeed in three places with 55 lines of code: (1) before the start and at the end of an epoch for Type-A bubbles, (2) after all FP operations preceding the first BP operation for Type-B bubbles, and (3) after the first BP operation following the last FP operation for Type-C bubbles. The instrumented code reports bubbles to the side task manager in FreeRide, as shown in step ➎ of Figure 3. The modifications are done once as the framework can be used for training different models.
@@
Figure 5. FreeRide System Architecture.
@@ 5. Use of Side Tasks Interface
Iterative programming interface. Figure 6 is an example of implementing a side task to train ResNet18 using the iterative interface of FreeRide in Python. Less important lines such as importing dependencies and parsing arguments are simplified. Porting this example involves mainly 6 steps. Step ➀: import FreeRide dependencies and inherit the iterative interface class, which includes an implementation for the state machine abstraction, communication with the side task manager, and the program-directed mechanism to limit the GPU execution time. The programmer only has to migrate the implementation of the original GPU workload to the interface. Steps ➁ and ➂: implement the side task initialization in 2 state transition functions, CreateSideTask() and InitSideTask(), to load the context into main memory and GPU memory respectively. Step ➃: implement the loop condition which checks if the step-wise execution is finished. Step ➄: implement the side task step in RunNextStep(). Step ➅: the main function handles argument parsing and runs the side task interface.
Most of the modifications are trivial, e.g., adding self. to variables such as model and optim in steps ➁, ➂ and ➄, which are required by Python. Aside from this, the programmer can directly copy the important logic, e.g., loading the dataset and training the model, from the original implementation. In addition, if the programmer customizes the model architecture instead of using the publicly available ones, the model implementation also does not require modification.
@@ 6.1. Methodology
We implement 3 types of side tasks: model training, graph analytics, and image processing using both iterative and imperative interfaces of FreeRide. Model training side tasks include ResNet18, ResNet50, and VGG19. We use the out-of-the-box models from PyTorch (pytorch_developers_models_nodate, ) and implement the training procedure ourselves. Graph analytics side tasks are adapted from Gardenia (xu_gardenia_2019, ). It includes PageRank (PR) which is based on the PageRank algorithm (page_pagerank_1999, ) and Graph SGD (SGD) which uses stochastic gradient descent to solve matrix factorization (koren_matrix_2009, ), both using the Orkut dataset (yang_defining_2012, ). The image processing (Image) side task resizes an input image and adds a watermark, which we adapt from Nvidia’s example (nvidia_developers_image_2019, ).
@@
We use the time increase $\Delta{T}$ and cost savings $S$ in Dollars due to side tasks as metrics. Time increase describes the performance overhead of co-locating side tasks with the main pipeline training workload. It is defined as $\Delta{T}=(T_{\rm withSideTasks}-T_{\rm noSideTask})/T_{\rm noSideTask}$, the ratio of extra time of pipeline training with side tasks, compared with the original DeepSpeed without any side tasks.
Cost savings describe the benefits of running side tasks. Since we cannot directly compare the throughput of different side tasks and the main pipeline training workload, we use their cost (dollars spent on GPUs) as a proxy. First, we define the cost of pipeline training without side tasks as
@@
and the extra cost of pipeline training due to side tasks as
$C_{\rm extra}=\Delta{T}\times C_{\rm noSideTask}~{}~{}~{}~{}.$
Then, we compute the cost of side tasks as if each of them were executed on a dedicated lower-tier GPU as
@@
$S=\frac{C_{\rm sideTasks}-C_{\rm extra}}{C_{\rm noSideTask}}~{}~{}~{}.$
@@
Table 1. Time increase $\Delta{T}$ (lower the better) and cost savings $S$ (positive=benefit, negative=loss, higher the better) of running DeepSpeed with side tasks.
@@
Iterative Imperative MPS Naive
Side task $\Delta{T}$ ${S}$ $\Delta{T}$ ${S}$ $\Delta{T}$ ${S}$ $\Delta{T}$ ${S}$
@@
6.2. Performance Evaluation
We run DeepSpeed to train a 3.6B model for 128 epochs with side tasks from Section 6.1.4 and compare the performance overhead, i.e., time increase ($\Delta{T}$) and cost savings ($S$) of using FreeRide with the two interfaces and the two comparative methods (as mentioned in Section 6.1.2). For model training side tasks, we set the batch size to 64. We run the same side task in all workers if they have enough GPU memory. We also run a mixed workload with 4 side tasks: PageRank, ResNet18, Image, and VGG19, each in one worker corresponding to the GPU of stages 0—3 in Section 6.1.3, respectively.
The results are summarized in Table 1. FreeRide consistently exhibits lower overhead than the comparative methods, showing only a 1.1% average time increase while achieving 7.8% average cost savings through side tasks using the iterative interface. The imperative interface achieves comparable cost savings but with a higher overhead as it relies on the less efficient framework-enforced mechanism to limit the side task’s execution time (Section 4.5). In comparison, the average time increase and cost savings for MPS are 48.7% and -4.5%, and for Naive are 54.7% and -29.2%. Their negative cost savings indicate that these approaches can increase the total cost. Notably, the time increase of Graph SGD with MPS is as high as 231.0%. This anomaly is due to Graph SGD’s high compute intensity. We conclude that FreeRide effectively utilizes bubbles in pipeline training for serving side tasks. While the comparative methods can utilize bubbles, unlike FreeRide, they are not designed for this purpose. Thus, they are inefficient in using bubbles, leading to higher costs.
@@ 6.3. Sensitivity Study
We change the side task batch size, DeepSpeed model size, and DeepSpeed micro-batch numbers of different side tasks, and study the time increase and cost savings of FreeRide with the iterative interface.
@@ 7. Related Work
Pipeline parallelism and bubbles. Prior research has aimed to improve the schedule of pipeline training to reduce bubbles (rasley_deepspeed_2020, ; deepspeed_developers_pipeline_2023, ; kim_torchgpipe_2020, ; huang_gpipe_2019, ; fan_dapple_2021, ; liu_hanayo_2023, ; li_chimera_2021, ; harlap_pipedream_2018, ; narayanan_memory-efficient_2021, ; tang_adaptive_2024, ; qi_zero_2023, ), and to leverage bubbles in pipeline parallelism by assigning specialized procedures coupled with pipeline training to gain benefits (osawa_pipefisher_2023, ; thorpe_bamboo_2023, ), as discussed in Section 2. In contrast, FreeRide does not require any changes to, or coupling with, pipeline training to serve generic GPU side tasks.
GPU sharing. Previous work has enhanced GPU sharing through optimized GPU primitives (yu_salus_2019, ), better scheduling of containers or VMs (zhao_efficient_2018, ; wu_transparent_2023, ), and tailored applications like training frameworks or model compilers (xiao_antman_2020, ; xiao_gandiva_2018, ; zhang_pilotfish_2022, ; bai_pipeswitch_2020, ; lim_2021_zico, ; han_2022_reef, ; veltair, ). Orion supports GPU sharing by intercepting and scheduling the CUDA kernel calls made by PyTorch(strati_orion_2024, ). However, Orion is agnostic of the bubbles in pipeline parallelism and can cause significant performance overhead, while FreeRide manages side tasks based on bubbles and achieves an extremely low performance overhead of about 1%. In addition, Orion only supports deep learning workloads implemented with PyTorch while FreeRide supports generic GPU side tasks. MPS and MIG (nvidia_developers_multi-process_2024, ; nvidia_developers_mig, ) are mechanisms provided by Nvidia for GPU sharing and virtualization. FreeRide leverages MPS to impose GPU memory limits on side tasks.

### v2

@@ 4.3. Profiling Bubbles and Side Tasks
Bubbles. To know the shapes of bubbles before serving side tasks with them, FreeRide runs DeepSpeed, monitors its estimated SM occupancy and GPU memory consumption through the PyTorch profiler (pytorch_profiler, ), and automatically measures each bubble’s duration and available GPU memory. Since the pipeline schedule determines bubbles, this offline profiling is done only once for each model and pipeline scheduling on the same hardware platform.
Side tasks. After the programmer has implemented the side task, FreeRide profiles it with the automated profiling tool for its performance characteristics of GPU memory consumption and per-step duration, which FreeRide uses for side task management and limiting GPU resources. For side tasks implemented using the iterative interface, this procedure is fully automated. The profiling tool runs the side task, monitors its GPU memory consumption, and records the timestamps at the start and end of RunNextStep() state transition for the per-step duration. For side tasks implemented using the imperative interface, the tool profiles GPU memory consumption in the same way. However, since the side task is not step-wise, the automated profiling tool does not measure the per-step duration.
@@ 4.4. Side Task Management
FreeRide’s side task management has two main roles. First, upon receiving a new side task, the side task manager assigns it to a suitable side task worker. Second, when the pipeline training system adds bubbles to the side task manager, the side task manager initiates the state transitions of side tasks (Figure 4(a)) through RPCs. In this way, the side tasks are only served during bubbles and do not compete for GPU resources with the main pipeline training workload.
@@
1:  Input: new task $Task$, workers’ metadata $Workers$
2:  $MinNumTasks\leftarrow\infty$
3:  $SelectedWorker\leftarrow None$
4:  for all $Worker$ in $Workers$ do
5:     if $Worker.GPUMem>Task.GPUMem$ then
6:        $NumTasks\leftarrow Worker.GetTaskNum()$
7:        if $NumTasks<MinNumTasks$ then
8:           $MinNumTasks\leftarrow NumTasks$
9:           $SelectedWorker\leftarrow Worker$
10:  if $SelectedWorker\neq None$ then
11:     $SelectedWorker.Add(Task)$
12:  else
13:     $RejectSideTask()$
Algorithm 1 describes how the side task manager assigns side tasks to workers. When the side task manager receives a new side task together with its GPU memory requirement (through profiling, Section 4.3), it first filters out all workers with enough available GPU memory (lines 4—5). Then, from these workers, it selects the one with the smallest number of tasks (lines 6—9). If the side task manager has selected a worker, it will assign the side task to that worker (lines 10—11). Otherwise, it will reject the side task because of insufficient GPU memory (line 13).
@@
1:  Input: workers’ metadata $Workers$
2:  while $SideTaskManagerIsRunning$ do
3:     for all $Worker$ in $Workers$ do
4:        if $Worker.CurrentBubble\neq None$ then
5:           if $Worker.CurrentBubble.HasEnded()$ then
6:              if $Worker.CurrentTask\neq None$ then
7:                 $Worker.CurrentTask.PauseSideTask()$
8:              $Worker.CurrentBubble\leftarrow None$
9:        if $Worker.HasNewBubble()$ then
10:           $Worker.UpdateCurrentBubble()$
11:           if $Worker.CurrentTask=None$ then
12:              if $Worker.TaskQueue.IsEmpty()$ then
13:                 $continue$
14:              $nextTask\leftarrow Worker.TaskQueue.Next()$
15:              $Worker.CurrentTask\leftarrow nextTask$
16:           if $Worker.CurrentTask.IsCreated()$ then
17:              $Worker.CurrentTask.InitSideTask()$
18:           else if $Worker.CurrentTask.IsPaused()$ then
19:              $Worker.CurrentTask.StartSideTask()$
Algorithm 2 describes how the side task manager manages bubbles and side tasks during pipeline training. The side task manager iterates through all workers (line 3). If CurrentBubble has just ended for a worker, the side task manager will pause CurrentTask of the worker and clear CurrentBubble (lines 4—8). Upon a new bubble, the side task manager updates the CurrentBubble of this worker (lines 9—10). It then checks if the worker has a CurrentTask. If not, it will select the one with the smallest submission timestamp from TaskQueue as CurrentTask (lines 11–15). After that, the side task manager initiates InitSideTask() if the newly added CurrentTask is in CREATED state (lines 16—17); otherwise, its state is PAUSED and the side task manager initiates StartSideTask() (lines 18—19).
@@ 4.5. GPU Resource Limit
GPU Memory. FreeRide leverages MPS to impose GPU memory limit on side tasks. I.e., when a worker creates a side task, it sets GPU memory limits using MPS. The side task process triggers an out-of-memory (OOM) error when its memory consumption exceeds the limit, but other processes remain unaffected. However, FreeRide is also compatible with other mechanisms for limiting GPU memory, e.g., multi-instance GPU (MIG) (nvidia_developers_mig, ) or manually implemented accounting through intercepting CUDA kernel calls (strati_orion_2024, ).
@@ 4.6. Implementation
We use DeepSpeed 0.12.2 (deepspeed_0_12_2, ) as the framework for pipeline training. We modify DeepSpeed in three places with 55 lines of code: (1) before the start and at the end of an epoch for Type-A bubbles, (2) after all FP operations preceding the first BP operation for Type-B bubbles, and (3) after the first BP operation following the last FP operation for Type-C bubbles. The instrumented code reports bubbles to the side task manager in FreeRide, as shown in step ➎ of Figure 3. The modifications are done once, as the framework can be used for training different models.
@@
Figure 5. Architecture of FreeRide.
@@ 5. Use of Side Tasks Interface
Iterative programming interface. Figure 6 is an example of implementing a side task to train ResNet18 using the iterative interface of FreeRide in Python. Less important lines such as importing dependencies and parsing arguments are simplified. Porting this example mainly involves 6 steps. Step ➀: import FreeRide dependencies and inherit the iterative interface class, which includes an implementation for the state machine abstraction, communication with the side task manager, and the program-directed mechanism to limit the GPU execution time. The programmer only needs to adapt the original GPU workload to the interface. Steps ➁ and ➂: implement the side task initialization in 2 state transition functions, CreateSideTask() and InitSideTask(), to load the context into main memory and GPU memory respectively. Step ➃: wrap the original loop implementation with RunNextStep(). Step ➄: the main function parses arguments and runs the side task interface.
Most modifications are trivial, e.g., wrapping implementations with side task state transition functions in Step ➁, ➂, and ➃, which are required by Python. Aside from this, the programmer can directly copy the important logic, e.g., loading the dataset and training the model, from the original implementation. In addition, if the programmer customizes the model architecture, the model implementation does not require modification.
@@ 6.1. Methodology
We implement 3 types of side tasks: model training, graph analytics, and image processing using both the iterative and the imperative interfaces of FreeRide. Model training side tasks include ResNet18, ResNet50, and VGG19. We implement the training procedure using out-of-the-box models from PyTorch (pytorch_developers_models_nodate, ). Graph analytics side tasks are adapted from Gardenia (xu_gardenia_2019, ). It includes PageRank (PR) which is based on the PageRank algorithm (page_pagerank_1999, ) and Graph SGD (SGD) which uses stochastic gradient descent to solve matrix factorization (koren_matrix_2009, ), both using the Orkut dataset (yang_defining_2012, ). The image processing (Image) side task resizes an input image and adds a watermark, which we adapt from Nvidia’s code (nvidia_developers_image_2019, ).
@@
We use time increase $I$ and cost savings $S$ in Dollars due to side tasks as metrics. Time increase describes the performance overhead of co-locating side tasks with the main pipeline training workload. It is the ratio of extra time of pipeline training with side tasks, compared with the original DeepSpeed without any side tasks, and lower time increase means lower overhead. It is defined as
$I=\frac{T_{\rm withSideTasks}-T_{\rm noSideTask}}{T_{\rm noSideTask}}~{}~{}~{}.$
Cost savings describe the benefits of running side tasks. It is hard to directly measure the benefits of running side tasks for two reasons. First, the throughput of different size tasks and the main pipeline training workload cannot be directly compared. Second, the workloads of side tasks are typically deployed on GPUs of smaller scales, while pipeline training mostly uses flagship GPUs. To compare the value of side tasks and pipeline training that runs on different GPUs with different throughputs and to calculate the benefits, we use their costs (dollars spent on GPUs) as a proxy. First, we define the cost of pipeline training without side tasks as
@@
and the cost of pipeline training with side tasks as
$C_{\rm withSideTasks}=P_{\rm Server-I}\times T_{\rm withSideTasks}~{}~{}~{}.$
Then, we compute the cost of running the same side tasks on dedicated lower-tier GPUs as
@@
$S=\frac{C_{\rm sideTasks}-(C_{\rm withSideTasks}-C_{\rm noSideTask})}{C_{\rm noSideTask}}~{}~{}~{}.$
$C_{\rm sideTasks}$ considers the cost of the side tasks served during bubbles, measured by the cost of running the same side task workloads on dedicated GPUs. $C_{\rm withSideTasks}-C_{\rm noSideTask}$ is the extra cost from co-location, measured by the increased costs due to co-locating side tasks and pipeline training.
6.2. Performance Evaluation
@@
We run DeepSpeed to train a 3.6B model for 128 epochs with side tasks from Section 6.1.4 and compare the performance overhead, i.e., time increase ($I$) and cost savings ($S$) of using FreeRide with the two interfaces and the two comparative methods (as mentioned in Section 6.1.2). For model training side tasks, we set the batch size as 64. We run the same side task in all workers if they have enough GPU memory. We also run a mixed workload with 4 side tasks: PageRank, ResNet18, Image, and VGG19, each in one worker corresponding to the GPU of stages 0—3 in Section 6.1.3, respectively.
Table 1. Throughput of GPU side tasks on different platforms measured as iterations per second.
Side task Iterative Server-II Server-CPU
ResNet18 1586.6 998.7 26.5
ResNet50 533.1 393.4 9.1
VGG19 170.7 161.8 3.0
PageRank 333.9 126.3 11.1
Graph SGD 4.2 1.5 0.6
Image 12.2 7.8 1.6
Performance compared to lower-tier GPU and CPU. Table 1 compares the throughput of side task workloads in Section 6.1.4 running on bubbles using the iterative interface of FreeRide (the Iterative column), on Server-II and Server-CPU as introduced in Section 6.1. The throughput is measured as iterations per second. For ResNet18, ResNet50, and VGG19 GPU workloads, one iteration corresponds to one batch of training data. For PageRank and Graph SGD, in each iteration, the graph algorithm runs over the input graph for one step. For image processing workload, one iteration processes one image.
This comparison shows that although side tasks run only during bubbles, they still achieve higher throughput compared to running on a standalone lower-tier GPU (RTX 3080) or the 8-core CPU instance. By introducing little overhead to pipeline parallel training, FreeRide harvests GPU resources that support a throughput of 1.06–2.82$\times$ of a standalone lower-tier GPU, and 7–59.9$\times$ of the CPU. These results demonstrate FreeRide’s effectiveness in harvesting the bubbles in pipeline parallel training.
Table 2. Time increase $I$ (lower the better) and cost savings $S$ (positive=benefit, negative=loss, higher the better) of running DeepSpeed with side tasks.
@@
Iterative Imperative Nvidia MPS Naive co-location
Side task $I$ ${S}$ $I$ ${S}$ $I$ ${S}$ $I$ ${S}$
@@
Co-location performance. Table 2 shows the time increase and cost savings of running DeepSpeed with side tasks of different co-location methods. FreeRide consistently exhibits lower overhead than the comparative methods, showing only a 1.1% average time increase while achieving 7.8% average cost savings through side tasks using the iterative interface. The imperative interface achieves comparable cost savings but with a higher overhead as it relies on the less efficient framework-enforced mechanism to limit the side task’s execution time (Section 4.5). In comparison, the average time increase and cost savings for MPS are 48.7% and -4.5%, and for Naive are 54.7% and -29.2%. Their negative cost savings indicate that these approaches can increase the total cost. Notably, the time increase of Graph SGD with MPS is as high as 231.0%. This anomaly is due to Graph SGD’s high compute intensity. We conclude that FreeRide effectively utilizes bubbles in pipeline training for serving side tasks. While the comparative methods can utilize bubbles, unlike FreeRide, they are not designed for this purpose. Thus, they are inefficient in using bubbles, leading to higher costs.
@@ 6.3. Sensitivity Study
We change the side task batch size, DeepSpeed model size, and DeepSpeed micro-batch numbers of side tasks, and study the time increase and cost savings of FreeRide with the iterative interface.
@@ 7. Related Work
Pipeline parallelism and bubbles. Prior work has aimed to improve the schedule of pipeline training to reduce bubbles (rasley_deepspeed_2020, ; deepspeed_developers_pipeline_2023, ; kim_torchgpipe_2020, ; huang_gpipe_2019, ; fan_dapple_2021, ; liu_hanayo_2023, ; li_chimera_2021, ; harlap_pipedream_2018, ; narayanan_memory-efficient_2021, ; tang_adaptive_2024, ; qi_zero_2023, ). Other work leverages bubbles in pipeline parallelism by assigning specialized procedures coupled with pipeline training to enable fault tolerance through replicated computation (thorpe_bamboo_2023, ), or for accelerating the training algorithms (osawa_pipefisher_2023, ; hydro, ). These approaches require changes to the training framework and are limited to certain types of workloads. In contrast, FreeRide does not require any changes to, or coupling with, pipeline training to serve generic GPU side tasks.
GPU sharing. Gandiva time-slices GPUs for multiple jobs with fallback to non-sharing GPUs (xiao_gandiva_2018, ). Salus designs job switching and memory sharing primitives for GPU sharing (yu_salus_2019, ). PipeSwitch further improves GPU sharing by designing fast context switch mechanisms between the host memory and the GPU memory (bai_pipeswitch_2020, ). AntMan designs dynamic scaling mechanisms for distributed deep learning workloads (xiao_antman_2020, ). Zico tracks the GPU memory allocation and reclamation of deep learning jobs and shares the GPU memory reclaimed by one job with other jobs (lim_2021_zico, ). Veltair proposes a compiler that co-optimizes the compiling results of co-located GPU workloads (veltair, ). TGS achieves GPU sharing of deep learning workloads in container clouds through rate control and transparent shared memory at the OS level (wu_transparent_2023, ). Recently, Orion proposes GPU sharing by intercepting and scheduling CUDA kernel calls made by PyTorch (strati_orion_2024, ). While these approaches propose methods to share GPUs and continuously improve the utilization of GPUs, they share the GPUs aggressively, without minimizing the impact on high-priority and high-cost workloads. Therefore, they would cause high overhead (time increase) if used to co-locate LLM training and side tasks, and subsequently yield little to no cost savings due to their high overhead. In comparison, FreeRide achieves GPU sharing while maintaining a very low overhead. PilotFish on the other hand leverages the free cycles in cloud gaming for deep learning workloads (zhang_pilotfish_2022, ), while FreeRide harvests the bubbles in pipeline parallelism for other generic GPU workloads. MPS and MIG (nvidia_developers_multi-process_2024, ; nvidia_developers_mig, ) are mechanisms provided by Nvidia for GPU sharing and virtualization. FreeRide leverages MPS to impose GPU memory limits on side tasks.